a)Briefly explain 3 principle of insurance

b) what are the advantages of discounting in finance

c) you have to make a payment on a load with a current balance of kshs 100,000 that matures 115days from now , interest accumulates daily on this loan at a rate of 6% per annum

I) what is the effective annual rate on this loan .

II) What error do you make in your calculations if you assume that interest is compounded continuously.



1
Expert's answer
2021-01-26T03:32:07-0500

a)

1. the principle of having an insurance interest;

2. the principle of insurability of risk;

3. the principle of equivalence.


b)The advantage of this method is due to the fact that cash flow management plays a huge role for all parties interested in the effective operation of the company, allows you to manage the value of the existing business and allows you to increase the financial flexibility of the company. Cash flow, in contrast to the net profit indicator, allows you to correlate cash inflows and outflows, taking into account depreciation and amortization, capital investments, receivables, changes in the structure of the company's own working capital.


c)

"FV=PV(1+ir)=100000(1+0.06\frac{115}{365})=101890.41"


i)"re=(1+\frac{i}{n})^n-1"

"re=(1+\frac{0.06}{365})^{365}-1=0.0618" or 6.18%


ii)Simple interest refers to the interest that is accrued only on the initial amount for each specific period of time.

Compound interest is a slightly different form of interest accrual. The interest here is not charged on the initial payment, but on the whole amount, together with the already accrued interest, which is currently in the account. That is, at the end of each period, the amount for which interest is accrued increases proportionally.

With a simple system, the profit increases in arithmetic progression, and with a complex system in geometric progression.